OpenStudy (piercetheveil47):

help? solve for s . s over 9 minus 23 is less than or equal to 37 A.) 2 B.) 7 C.) 126 D.) 540

OpenStudy (texaschic101):

(s/9) - 23 < = 37 s/9 <= 37 + 23 s/9 <= 60 -- multiply both sides by 9 s <= 540

OpenStudy (piercetheveil47):

thank you so much can you help me with one more?? @texaschic101

OpenStudy (texaschic101):

possibly...what ya got ?

OpenStudy (piercetheveil47):

@texaschic101 solve for p. 6p-19 < 29. A.) P < 2 B.) P<8 C.) P <12 D.) P <288 Thanks so much that was my last one :)

OpenStudy (texaschic101):

6p - 19 < 29 -- add 19 to both sides 6p < 29 + 19 6p < 48 -- divide both sides by 6 p < 48/6 p < 8

OpenStudy (piercetheveil47):

thank you oh my godddd :,) <3 i got a 100 dude thank you so much @texaschic101

OpenStudy (texaschic101):

happy to help :)